# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v4

@@ Appendix A Detailed Derivation on Sorting Boundary Zero Modes of Toeplitz $K$-matrices
from which we know that $\beta_{1}\beta_{2}=1$. $\mathbf{u}$ is required to localize at one $z$-boundary, i.e., all the $\{\beta_{j}\}$ in one trial solution $\mathbf{u}$ satisfies either $|\beta_{j}|>1$ or $|\beta_{j}|<1$. If a trial solution $\mathbf{u}$ is a linear combination of boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|<1$, it is required to match the boundary condition
$A\begin{pmatrix}u_{1}^{a}\\ u_{1}^{b}\end{pmatrix}+B^{T}\begin{pmatrix}u_{2}^{a}\\ u_{2}^{b}\end{pmatrix}=0.$ (94)
If a trial solution $\mathbf{u}$ is a linear combination of boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|>1$, it is required to match the boundary condition
$B\begin{pmatrix}u_{N-1}^{a}\\ u_{N-1}^{b}\end{pmatrix}+A\begin{pmatrix}u_{N}^{a}\\ u_{N}^{b}\end{pmatrix}=0.$ (95)
As is discussed in the main text, we classify our discussion based on the order of Eq. (88), which affects the # of boundary bases. A detailed discussion on $\det B=0$ cases are presented in Sec. A.1, while a detailed discussion on $\det B\neq 0$ cases are presented in Sec. A.2.
A.1 $\det B$=0
Discussion on $\mathbf{w}^{1}$ satisfying the boundary condition Eq. (94)
If $l_{4}\neq 0$, for the boundary base $\mathbf{w}^{1}$ with entries
@@ A.1 $\det B$=0
where $c_{1}\neq 0$ is a normalization constant. Since $10<|\beta_{1}|<$ 1, $\mathbf{w}^{1}$ is required to satisfy the boundary condition Eq. (94), which renders
$\beta_{1}\left(\beta_{1}\left(\beta_{1}^{2}+1\right)l_{1}l_{4}-m(\beta_{1}(\beta_{1}l_{2}+m)+l_{3})-\beta_{1}l_{3}(\beta_{1}(\beta_{1}l_{2}+m)+l_{3})+\left(\beta_{1}^{2}+1\right)l_{4}n\right)=0,$ (97a)
@@
If $l_{4}\neq 0$, Eq. (97b) implies $\beta_{1}(l_{2}-l_{3})+m=0$, rendering either $\left\{\begin{array}[]{l}l_{2}=l_{3}\\ m=0\end{array}\right.$ or $\beta_{1}=\frac{m}{l_{3}-l_{2}}$. $\left\{\begin{array}[]{l}l_{2}=l_{3}\\ m=0\end{array}\right.$ together with Eq. (97a) enforces $|\beta_{1}|=1$ or $|\beta_{1}|=0$, which does not correspond to a boundary zero mode and should be discarded. Hence we shall continue with $\beta_{1}=\frac{m}{l_{3}-l_{2}}$. Since $l_{4}\neq 0$, we can safely set $l_{1}=l_{2}l_{3}/l_{4}$. Eq. (97a) and Eq. (93) render
@@
Therefore, if $02l_{3}m-l_{4}n=$ 0, Eq. (96) represents a boundary state that satisfies the boundary condition Eq. (94). Taking the aforementioned conditions into account, if
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.$ (101)
@@
The above discussion ignored $l_{4}=0$. If $l_{4}=0$, $\det B=0$ necessitates $l_{2}l_{3}=0$. If $l_{2}=l_{3}=l_{4}=0$, i.e.
@@
the bulk recurrence relation Eq. (86) and the boundary condition Eq. (94) render
@@
Nontrivial solution to Eq. (102) enforces $m=0$, resulting in a zero determinant for the $K$-matrix. Therefore, $l_{2}=l_{3}=l_{4}=0$ is excluded from consideration. If $l_{2}=0$, $l_{3}\neq 0$, i.e.
@@
where $c_{1}\neq 0$ is the normalization constant. Since $10<|\beta_{1}|<$ 1, $\mathbf{w}^{1}$ is required to satisfy the boundary condition Eq. (94), which gives
@@
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{4}=l_{2}=0,\\ |m|<|l_{3}|,\end{array}\right.$ (108)
there exists a boundary zero mode $\mathbf{u}^{1}=\mathbf{w}^{1}$ given by Eq. (96), satisfying boundary condition Eq. (94), where
$w^{1,a}=0,\quad w^{1,b}=\frac{1}{\beta_{1}}\sqrt{\frac{1-\beta_{1}^{2N}}{1-\beta_{1}^{2}}},\quad\beta_{1}=-\frac{m}{l_{3}}.$
@@
$\left\{\begin{array}[]{l}w^{1,a}=0\\ w^{1,b}=c_{1}\\ \beta_{1}=-\frac{l_{2}}{m}\end{array}\right.\quad\text{or}\quad\left\{\begin{array}[]{l}w^{1,a}=-(l_{2}+m\beta_{1})c_{1}\\ w^{1,b}=(l_{1}+n\beta_{1}+l_{1}\beta_{1}^{2})c_{1}\\ \beta_{1}=-\frac{m}{l_{2}}\end{array}\right.,$ (115)
where $c_{1}\neq 0$ is a normalization constant. In this condition, only if $n=l_{1}=l_{3}=l_{4}=0$, $|m|<|l_{2}|$, i.e.
$A=\begin{pmatrix}0&m\\ m&0\end{pmatrix},\quad B=\begin{pmatrix}0&l_{2}\\ 0&0\end{pmatrix},$ (116)
there is a boundary zero mode $\mathbf{u}^{1}=\mathbf{w}^{1}$ given by Eq. (96), satisfying the boundary condition Eq. (94), where
$w^{1,a}=\frac{1}{\beta_{1}}\sqrt{\frac{1-\beta_{1}^{2N}}{1-\beta_{1}^{2}}},\quad w^{1,b}=0,\quad\beta_{1}=-\frac{m}{l_{2}}.$
Discussion on $\mathbf{w}^{2}$ satisfying the boundary condition Eq. (95)
@@
where $c_{2}\neq 0$ is a normalization constant. Since $|\beta_{2}|>1$, it is required to satisfy the boundary condition Eq. (95), rendering
$-\left(\beta_{2}^{2}+1\right)l_{4}(l_{1}+n\beta_{2})+\beta_{2}^{2}l_{2}^{2}+l_{2}\left(l_{3}+\beta_{2}m+\beta_{2}^{3}m\right)+\beta_{2}^{2}m(l_{3}+\beta_{2}m)=0,$ (118a)
@@
Therefore, if $02l_{2}m-l_{4}n=$ 0, Eq. (117) represents a boundary zero mode that satisfies the boundary condition Eq. (95). Taking the aforementioned conditions into account, if
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.$ (122)
@@
the bulk recurrence relation Eq. (86) and the boundary condition Eq. (95) render
@@
Nontrivial solution to Eq. (123) enforces $m=0$, resulting in a zero determinant for the $K$-matrix. Therefore, $l_{2}=l_{3}=l_{4}=0$ is excluded from consideration. If $l_{3}=0$, $l_{2}\neq 0$, i.e.
@@
$w^{2,a}=0,\quad w^{2,b}=c_{2},$ (124)
where $c_{2}\neq 0$ is a normalization constant. Since $|\beta_{2}|>1$, $\mathbf{w}^{2}$ is required to satisfy the boundary condition Eq. (95), which gives
@@
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{4}=l_{3}=0,\\ |m|<|l_{2}|,\end{array}\right.$ (129)
@@
$w^{2,a}=0,\quad w^{2,b}=\frac{1}{\beta_{2}}\sqrt{\frac{1-\beta_{2}^{2N}}{1-\beta_{2}^{2}}},\quad\beta_{2}=-\frac{l_{2}}{m},$
satisfying the boundary condition Eq. (95). If $l_{2}=0$, $l_{3}\neq 0$ i.e.
@@
$\left\{\begin{array}[]{l}w^{2,a}=-m\beta_{2}c_{2}\\ w^{2,b}=(l_{1}+n\beta_{2})c_{2}\\ \beta_{2}=-\frac{l_{3}}{m}\end{array}\right.,$ (133)

### v5

@@ Appendix A Detailed Derivation on Sorting Boundary Zero Modes of Toeplitz $K$ matrices
from which $\beta_{1}\beta_{2}=1$ is immediately apparent. Here we require $\beta_{1}$ satisfies $10<|\beta_{1}|<$ 1, and the corresponding boundary base is denoted as $\mathbf{w}^{1}$. The other solution $\beta_{2}$ satisfies $|\beta_{2}|>1$, of which the corresponding boundary base is denoted as $\mathbf{w}^{2}$. In Appendix B, we provide detailed discussions on the conditions for the elements of the $K$ matrix that enable trial solutions $\mathbf{u}^{1}=\mathbf{w}^{1}$ or $\mathbf{u}^{2}=\mathbf{w}^{2}$ being legitimate boundary zero modes. The discussion on $\mathbf{u}^{1}$ satisfying the lower boundary condition Eq. (151a) is presented in Subsection B.1, while the discussion on $\mathbf{u}^{2}$ satisfying the upper boundary condition Eq. (151b) is presented in Subsection B.2. Furthermore, the discussion in Subsection B.1 or B.2 is categorized into $l_{4}\neq 0$ and $l_{4}=0$, which is due to the explicit form of $\mathbf{w}^{1,2}$ being different in these cases. Finally, the condition for the appearance of boundary zero modes under the $\det B=0$ condition are summarized in Subsection B.3. We discover that the following set of conditions yields a boundary zero mode satisfying the boundary condition Eq. (151a):
$\mathbb{A}_{1}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{A}_{2}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{4}=l_{2}=0,\\ |m|<|l_{3}|,\end{array}\right.\quad\text{or}\quad\mathbb{A}_{3}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ nl_{2}=2l_{1}m,\quad l_{3}=l_{4}=0,\\ |m|<|l_{2}|.\end{array}\right.$ (162)
Similarly, the following set of conditions yields a boundary zero mode satisfying the boundary condition Eq. (151b):
$\mathbb{B}_{1}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{B}_{2}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{4}=l_{3}=0,\\ |m|<|l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{B}_{3}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ nl_{3}=2l_{1}m,\quad l_{2}=l_{4}=0,\\ |m|<|l_{3}|.\end{array}\right.$ (172)
If $\det B\neq 0$, Eq. (146) is a fourth-degree equation. In this case, if the element $l_{4}$ in block $B$ is nonzero, carrying on a comprehensive analytical analysis is challenging and requires a case-by-case study. To simplify discussion, we focus on the special case where $l_{4}=0$. In this situation, the equation for $\{\beta_{j}\}$ (Eq. (146)) factorizes into
$(l_{2}+m\beta_{j}+l_{3}\beta_{j}^{2})(l_{3}+m\beta_{j}+l_{2}\beta_{j}^{2})=0,\quad j=1,2,3,4,$ (173)
which facilitates our discussion. The detailed discussion is presented in Appendix C, where we divide our discussion into Subsection C.1 and Subsection C.2, addressing boundary conditions Eq. (151a) and Eq. (151b), respectively. Within each subsection, the discussion is organized into four distinct cases: Case I $\sim$ Case IV. For each case, trial solutions are constructed using the boundary bases. Each trial solution, denoted as $\mathbf{u}^{i,j}$, is a linear superposition of two boundary bases, $\mathbf{w}^{i}$ and $\mathbf{w}^{j}$ (Eq. (143)), where the moduli of parameters $\beta_{i}$ and $\beta_{j}$ are either both greater than 1 or both less than 1. These trial solutions are required to match the corresponding boundary conditions. This procedure yields explicit constraints on the elements of the $K$ matrix that guarantee the existence of boundary zero modes. We discover that the following conditions yield a boundary zero mode satisfying the boundary condition Eq. (151a):
$\mathbb{C}_{1}:\left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\\ \end{array}\right.\quad\text{or}\quad\mathbb{C}_{2}:\left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\end{array}\right.$ (180)
Similarly, the following conditions yield a boundary zero mode satisfying the boundary condition Eq. (151b):
$\mathbb{D}_{1}:\left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{D}_{2}:\left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\\ \end{array}\right.$ (187)
Taking the union of (162) and (180), we find that the $K$ matrix Eq. (2) has a boundary zero mode matching the boundary condition Eq. (151a) if one of the following conditions is satisfied:
$\mathbb{M}_{1}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{M}_{2}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\\ \end{array}\right.\quad\text{or}\quad\mathbb{M}_{3}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\end{array}\right.$ (197)
Similarly, taking the union of (172) and (187), we find that the $K$ matrix Eq. (2) has a boundary zero mode matching the boundary condition Eq. (151b) if one of the following conditions is satisfied:
$\mathbb{N}_{1}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{N}_{2}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|,\\ \end{array}\right.\quad\text{or}\quad\mathbb{N}_{3}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|.\end{array}\right.$ (207)
The condition for the $K$ matrix (Eq. (2)) possessing two boundary zero modes, i.e. being a Type-I $K$ matrix, is obtained by the set operation $(\mathbb{M}_{1}\cup\mathbb{M}_{2}\cup\mathbb{M}_{3})\cap(\mathbb{N}_{1}\cup\mathbb{N}_{2}\cup\mathbb{N}_{3})$. The result is shown in the first row of Table 1.
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq l_{3},\\ |m|<|l_{2}+l_{3}|,\\ l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m.\end{array}\right.$ (211)
Furthermore, set operation $(\mathbb{M}_{1}\cup\mathbb{M}_{2}\cup\mathbb{M}_{3}\cup\mathbb{N}_{1}\cup\mathbb{N}_{2}\cup\mathbb{N}_{3})\backslash[(\mathbb{M}_{1}\cup\mathbb{M}_{2}\cup\mathbb{M}_{3})\cap(\mathbb{N}_{1}\cup\mathbb{N}_{2}\cup\mathbb{N}_{3})]$ yields that the $K$ matrix Eq. (2) has only one boundary zero mode, i.e. a Type-II $K$ matrix, if the elements in blocks $A$ and $B$ satisfy one of the following conditions
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq l_{3},\\ |m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ |m|<|l_{2}-l_{3}|,\\ \det B=l_{1}l_{4}-l_{2}l_{3}=0,\\ (2l_{2}m-l_{4}n)(2l_{3}m-l_{4}n)=0.\end{array}\right.$ (220)
More specifically, $\left\{\begin{array}[]{l}l_{4}=0\\ |l_{2}|<|l_{3}|\end{array}\right.$ or $\left\{\begin{array}[]{l}l_{4}\neq 0\\ 2l_{3}m=l_{4}n\end{array}\right.$ indicates that the boundary zero mode is a lower boundary zero mode satisfying the boundary condition Eq. (151a); $\left\{\begin{array}[]{l}l_{4}=0\\ |l_{3}|<|l_{2}|\end{array}\right.$ or $\left\{\begin{array}[]{l}l_{4}\neq 0\\ 2l_{2}m=l_{4}n\end{array}\right.$ indicates that the boundary zero mode is an upper boundary zero mode satisfying the boundary condition Eq. (151b). The results are summarized in the second row of Table 1.
In order to streamline the lengthy discussion, a mind map is presented in Fig.LABEL:mindmap.
Appendix B The Conditions for the Existence of Boundary Zero Modes under $\det B$=0
Under the condition $\det B=0$, the equation for $\beta_{j}$ (Eq. (146)) reduces to a quadratic equation Eq. (152), from which $\beta_{1}\beta_{2}=1$ is immediately apparent. Here we require $\beta_{1}$ satisfies $10<|\beta_{1}|<$ 1, and the corresponding boundary base is denoted as $\mathbf{w}^{1}$. The other solution $\beta_{2}$ satisfies $|\beta_{2}|>1$, of which the corresponding boundary base is denoted as $\mathbf{w}^{2}$. Each trial solution consists of only one boundary base, i.e. $\mathbf{u}^{1}=\mathbf{w}^{1}$ and $\mathbf{u}^{2}=\mathbf{w}^{2}$. In this subsection, we are going to figure out the explicit conditions that enable boundary zero modes. The discussion on $\mathbf{u}^{1}$ and $\mathbf{u}^{2}$ are presented separately in Subsection B.1 and B.2. Furthermore, the discussions in Subsection B.1 and B.2 are categorized into $l_{4}\neq 0$ and $l_{4}=0$, which is due to the explicit form of $\mathbf{w}^{1,2}$ being different in these cases. The conditions for the appearance of boundary zero modes under $\det B=0$ are summarized in the final.
B.1 Discussion on $\mathbf{u}^{1}=\mathbf{w}^{1}$ satisfying the boundary condition Eq. (151a)
$l_{4}\neq 0$
For the boundary base $\mathbf{w}^{1}$ with entries
@@ Appendix B The Conditions for the Existence of Boundary Zero Modes under $\det B$=0
where $c_{1}\neq 0$ is a normalization constant. Since $10<|\beta_{1}|<$ 1, $\mathbf{w}^{1}$ is required to satisfy the boundary condition Eq. (151a), which renders
$\beta_{1}(\beta_{1}^{3}\left(l_{1}l_{4}-l_{2}l_{3}\right)+\beta_{1}\left(-l_{3}^{2}+l_{1}l_{4}-m^{2}\right)+\beta_{1}^{2}\left(-ml_{2}-l_{3}m+l_{4}n\right)-l_{3}m+l_{4}n)=0,$ (222a)
@@
Eq. (222b) implies $\beta_{1}(l_{2}-l_{3})+m=0$, rendering either $\left\{\begin{array}[]{l}l_{2}=l_{3}\\ m=0\end{array}\right.$ or $\beta_{1}=\frac{m}{l_{3}-l_{2}}$. $\left\{\begin{array}[]{l}l_{2}=l_{3}\\ m=0\end{array}\right.$ together with Eq. (222a) enforces $|\beta_{1}|=1$ or $|\beta_{1}|=0$, which does not correspond to a boundary zero mode and should be discarded. The physical meaning of $m$ also forbids $m=0$. Hence we shall continue with $\beta_{1}=\frac{m}{l_{3}-l_{2}}$. Since $l_{4}\neq 0$, we can safely set $l_{1}=l_{2}l_{3}/l_{4}$. Eq. (222a) and Eq. (152) render
@@
Therefore, if $02l_{3}m-l_{4}n=$ 0, Eq. (221) represents a boundary state that satisfies the boundary condition Eq. (151a). Taking the aforementioned conditions into account, we obtain that if
$\mathbb{A}_{1}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.$ (226)
@@
$l_{4}=0$
The above discussion ignored $l_{4}=0$. If $l_{4}=0$, $\det B=0$ necessitates $l_{2}l_{3}=0$.
If $l_{2}=l_{3}=l_{4}=0$, i.e.
@@
the bulk recurrence relation Eq. (144) and the boundary condition Eq. (151a) render
@@
Nontrivial solution to Eq. (227) enforces $m=0$, resulting in a zero determinant for the $K$ matrix. Therefore, $l_{2}=l_{3}=l_{4}=0$ is excluded from consideration.
If $l_{2}=0$, $l_{3}\neq 0$, i.e.
@@
where $c_{1}\neq 0$ is the normalization constant. Since $10<|\beta_{1}|<$ 1, $\mathbf{w}^{1}$ is required to satisfy the boundary condition Eq. (151a), which gives
@@
$\mathbb{A}_{2}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{4}=l_{2}=0,\\ |m|<|l_{3}|,\end{array}\right.$ (233)
there exists a boundary zero mode $\mathbf{u}^{1}=\mathbf{w}^{1}$ given by Eq. (221), satisfying the boundary condition Eq. (151a), where
$w^{1,a}=0,\quad w^{1,b}=\frac{1}{\beta_{1}}\sqrt{\frac{1-\beta_{1}^{2}}{1-\beta_{1}^{2N}}},\quad\beta_{1}=-\frac{m}{l_{3}}.$
@@
$w^{1,a}=-(l_{2}+m\beta_{1})c_{1},\quad w^{1,b}=(l_{1}+n\beta_{1}+l_{1}\beta_{1}^{2})c_{1},\quad\beta_{1}=-\frac{m}{l_{2}},$
where $c_{1}\neq 0$ is a normalization constant. Since $10<|\beta_{1}|<$ 1, $\mathbf{w}^{1}$ is required to satisfy the boundary condition Eq. (151a), which renders
$\begin{pmatrix}n\beta_{1}+l_{1}\beta_{1}^{2}&m\beta_{1}\\ m\beta_{1}+l_{2}\beta_{1}^{2}&0\end{pmatrix}\begin{pmatrix}w^{1,a}\\ w^{1,b}\end{pmatrix}=0.$
Therefore, if the elements of the $K$ matrix in Eq. (2) satisfy
$\mathbb{A}_{3}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ nl_{2}=2l_{1}m,\\ l_{3}=l_{4}=0,\\ |m|<|l_{2}|,\end{array}\right.$ (238)
then a boundary zero mode $\mathbf{u}^{1}=\mathbf{w}^{1}$ given by Eq. (221) exists that match the boundary condition Eq. (151a), where
$w^{1,a}=\frac{2\left(m^{2}-l_{2}^{2}\right)}{\beta_{1}\sqrt{(mn-2l_{1}l_{2})^{2}+4\left(l_{2}^{2}-m^{2}\right)^{2}}}\sqrt{\frac{1-\beta_{1}^{2}}{1-\beta_{1}^{2N}}},\ w^{1,b}=\frac{2l_{1}l_{2}-mn}{\beta_{1}\sqrt{(mn-2l_{1}l_{2})^{2}+4\left(l_{2}^{2}-m^{2}\right)^{2}}}\sqrt{\frac{1-\beta_{1}^{2}}{1-\beta_{1}^{2N}}},\ \beta_{1}=-\frac{m}{l_{2}}.$
B.2 Discussion on $\mathbf{u}^{2}=\mathbf{w}^{2}$ satisfying the boundary condition Eq. (151b)
$l_{4}\neq 0$
@@
where $c_{2}\neq 0$ is a normalization constant. Since $|\beta_{2}|>1$, it is required to satisfy the boundary condition Eq. (151b), rendering
$\beta_{2}^{3}\left(l_{2}m-l_{4}n+m^{2}\right)+\beta_{2}^{2}\left(l_{3}m+l_{2}^{2}-l_{1}l_{4}\right)+\beta_{2}\left(l_{2}m-l_{4}n\right)+l_{2}l_{3}-l_{1}l_{4}=0,$ (240a)
@@
Therefore, if $02l_{2}m-l_{4}n=$ 0, Eq. (239) represents a boundary zero mode that satisfies the boundary condition Eq. (151b). Taking the aforementioned conditions into account, if
$\mathbb{B}_{1}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.$ (244)
@@
the bulk recurrence relation Eq. (144) and the boundary condition Eq. (151b) render
@@
Nontrivial solution to Eq. (245) enforces $m=0$, resulting in a zero determinant for the $K$ matrix. Therefore, $l_{2}=l_{3}=l_{4}=0$ is excluded from consideration.
If $l_{3}=0$, $l_{2}\neq 0$, i.e.
@@
$w^{2,a}=0,\quad w^{2,b}=c_{2},\quad\beta_{2}=-\frac{l_{2}}{m},$ (246)
where $c_{2}\neq 0$ is a normalization constant. Since $|\beta_{2}|>1$, $\mathbf{w}^{2}$ is required to satisfy the boundary condition Eq. (151b), which gives
@@
$\mathbb{B}_{2}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{4}=l_{3}=0,\\ |m|<|l_{2}|,\end{array}\right.$ (251)
@@
$w^{2,a}=0,\quad w^{2,b}=\frac{1}{\beta_{2}}\sqrt{\frac{1-\beta_{2}^{2}}{1-\beta_{2}^{2N}}},\quad\beta_{2}=-\frac{l_{2}}{m},$
satisfying the boundary condition Eq. (151b).
If $l_{2}=0$, $l_{3}\neq 0$, i.e.
@@
$w^{2,a}=-(m+l_{3}\beta_{2})c_{2},\quad w^{2,b}=(l_{1}+n\beta_{2}+l_{1}\beta_{2}^{2})c_{2},\quad\beta_{2}=-\frac{l_{3}}{m},$
where $c_{2}\neq 0$ is a normalization constant. Since $|\beta_{2}|>1$, $\mathbf{w}^{2}$ is required to satisfy the boundary condition specified by Eq. (151b), which renders
$\left(\begin{array}[]{cc}l_{1}+\beta_{2}n&l_{3}+\beta_{2}m\\ \beta_{2}m&0\\ \end{array}\right)\begin{pmatrix}w^{2,a}\\ w^{2,b}\end{pmatrix}=0.$ (254)
Therefore, if the elements of the $K$ matrix in Eq. (2) satisfy
$\mathbb{B}_{3}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ nl_{3}=2l_{1}m,\\ l_{2}=l_{4}=0,\\ |m|<|l_{3}|,\end{array}\right.$ (259)